OpenStudy (anonymous):

Pls pls help - Suppose the velocity of a car starting from rest is given by V= 200t/(2t+8) where v is the velocity in mph and t is the time in seconds. Find the number of seconds when the velocity is 50 mph

OpenStudy (dumbcow):

solve for t V(2t+8) = 200t 2Vt + 8V = 200t 8V = 200t - 2Vt 8V/(200-2V) = t replace V with 50 t = 8*50/(200-(2*50)) t = 400/100 = 4